OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6.

OpenStudy (anonymous):

standard linear equation is y = mx + c gradient = -6 so eqn is y = -6x + c now to find c substitute the coordinates into the equation -3 = -6( -2) + c -3 = 12 + c c = -15 so y = -6x -15